# Reference-free Hallucination Detection for Large Vision-Language Models

### v1

@@ Abstract
Reference-free Hallucination Detection
for Large Vision-Language Models
${\textbf{Qing Li}}^{1}$, ${\textbf{Chenyang Lyu}}^{1}$, ${\textbf{Jiahui Geng}}^{1}$, ${\textbf{Derui Zhu}}^{2}$, ${\textbf{Maxim Panov}}^{1}$, ${\textbf{Fakhri Karray}}^{1}$ 1 Mohamed bin Zayed University of Artificial Intelligence {qing.li, chenyang.lyu, jiahui.geng, maxim.panov, fakhri.karray}@mbzuai.ac.ae 2 Technical University of Munich derui.zhu@tum.de
@@ 1 Introduction
Large vision-language models (LVLMs) like LLaVA Liu et al. (2023, 2024b), MiniGPT-4 Zhu et al. (2024b) have demonstrated remarkable capabilities in understanding and generating complex visual and textual content. However, an emerging concern with these models is their tendency towards hallucination Li et al. (2023); Zhou et al. (2024). For instance, a model might describe an object or event in the generated text that is not present in the input image. This phenomenon poses challenges for the reliability of LVLMs, highlighting the intrinsic limitations of current models in maintaining coherence between visual inputs and textual descriptions Huang et al. (2023a); Liu et al. (2024a).
Most existing methods of assessing hallucinations rely on external models. Li et al. (2023) proposed POPE, which requires automated segmentation tools like SEEM Zou et al. (2024) to identify present and absent objects in images. Faithscore Jing et al. (2023) utilizes a visual entailment model Xie et al. (2019) to validate the factuality, focusing on object attributes and their relationships. Nonetheless, these approaches increase the computational cost and are restricted by the capabilities of these models.
@@
Reference-free methods are a promising approach to address this issue, primarily utilizing the model’s own knowledge. They have attracted great research interest in Large Language Models (LLMs), with typical methods including the uncertainty-based methods Hu et al. (2023); Geng et al. (2024); Huang et al. (2023b); Vazhentsev et al. (2023); Fadeeva et al. (2023), consistency-based methods Manakul et al. (2023) and supervised uncertainty quantification (SUQ) approach Chen et al. (2023); Azaria and Mitchell (2023). However, their potential in the context of LVLMs is unknown. They have not yet been systematically studied, possibly due to the complexity of multimodal data. We are committed to bridging this gap.
In this work, we collect and implement a battery of state-of-the-art methods, including four metrics of uncertainty-based methods, four variants of consistency-based methods, and the supervised uncertainty quantification method. The experiments are conducted on five LVLMs with different types, versions, and sizes on Yes-and-No and Open-ended tasks. Our findings show that the SUQ method outperforms other approaches in hallucination detection, both at the sentence and passage levels.
The main contributions of this paper are summarized as follows:
@@ 2 Reference-free Hallucination Detection Methods
Uncertainty-based methods have the hypothesis that when LVLMs are uncertain about generated information, generated tokens often have higher uncertainty, shown in Figure 1(b). Uncertainty-based methods are extensively studied for hallucination detection in LLMs Hu et al. (2023); Geng et al. (2024); Huang et al. (2023b); Fadeeva et al. (2023, 2024), and are also used as benchmark techniques Zhu et al. (2024a); Manakul et al. (2023). We adopt the four proposed metrics: AvgProb, AvgEnt, MaxProb, MaxEnt to aggregate token-level uncertainty to measure sentence-level and passage-level uncertainty. Appendix A.1 provides detailed descriptions of these metrics.
@@ 3 Study Design
Models. We evaluate five representative categories of LVLMs: MiniGPT-4v Zhu et al. (2024b), LLaVA-v1.5-7b Liu et al. (2023), LLaVA-v1.6-vicuna-7b, LLaVA-v1.6-mistral-7b Liu et al. (2024b) and LLaVA-v1.6-vicuna-13b. These models, featuring different architectures and strong performance, are used as benchmarks in various works Huang et al. (2023a); Zhou et al. (2024). We only test models of a 7-billion and 13-billion size due to the limited computational resources.
4 Experimental Results and Analysis
The comprehensive results on Open-ended tasks are presented in Tables 1 and 2. The numbers in the pink and yellow background, respectively, represent the best results for the uncertainty-based methods and consistency-based methods, while the numbers in bold indicate the overall best results. We utilize AUC-PR Davis and Goadrich (2006), which stands for the area under the precision-recall curve, to objectively evaluate the effectiveness of each model. The higher the value of AUC-PR, the stronger the ability of this method for hallucination detection.
@@ 4.1 Effectiveness of Various Methods
The performance effects of the three types of methods on Yes-or-No tasks align with those on Open-ended tasks. Detailed results are demonstrated in Appendix C.1.
4.2 Robustness and Limitations of SUQ Methods.
@@
4.3 Impact of Image Clarity on Hallucination Detection Methods
@@
4.4 Impact of Data Source on various Methods
@@
Others. We have also explored the impact of unimportant tokens for uncertainty-based methods and the optimal hidden layers for SUQ. They are discussed respectively in Appendices C.2 and C.3.
@@ 5 Conclusion
This paper systematically studies reference-free approaches for detecting the hallucination of LVLMs. Our results indicate that the SUQ method performs best as a supervised method, but it relies on sufficient training data, which could limit its applicability in certain scenarios. Consistency-based methods outperform uncertainty-based methods; however, they require more computation and are more suitable for black-box models. Uncertainty-based methods are particularly well-suited for tasks similar to Yes-or-No questions. We call for more accurate and robust methods.
@@
Three major limitations are identified in this work.
First, we did not compare with reference-based methods. The performance of reference-based methods depends on the performance of external models and is specific to certain tasks. In this article, we primarily aim to explore the potential of reference-free methods, which have a broader range of applications.
Second, we did not classify types of hallucinations; limited by resources, even with fine-grained annotations, we only marked whether sentences contained hallucinations without distinguishing types such as object existence errors, attribute errors, text recognition errors, and fact-conflicting errors. This might be where SUQ methods have an advantage.
Third, we only studied white-box models. Recent research indicates that white-box models can be proxies for black-box models to calculate their confidence regarding specific issues. Unigram method is one such proxy model approach. In the future, we could use more complex models, closer in architecture to the target models, as proxies to explore their potential in hallucination detection.
@@ B.2 Experimental Setup
GPT-4v v1.5-7b v1.5-7b v1.6-Mistral
Adv Random 48.73 48.73 48.73 48.73
Uncertainty-based Methods
AvgProb 77.43 86.82 91.80 92.04
AvgEnt 48.89 50.05 48.90 48.92
Supervised Uncertainty Quantification Method
SUQ 87.23 92.35 93.32 93.50
Ran Baseline 48.73 48.73 48.73 48.73
Uncertainty-based Methods
AvgProb 86.62 95.89 96.23 96.99
AvgEnt 48.76 50.25 49.91 50.41
Supervised Uncertainty Quantification Method
SUQ 94.44 97.86 98.68 98.33
Pop Baseline 48.73 48.73 48.73 48.73
Uncertainty-based Methods
AvgProb 82.11 92.73 93.57 95.06
AvgEnt 48.50 49.76 50.63 51.39
Supervised Uncertainty Quantification Method
SUQ 91.02 97.85 98.11 98.50
GQA Baseline 48.47 48.47 48.47 48.47
Uncertainty-based Methods
AvgProb 66.87 78.45 80.31 79.15
AvgEnt 48.75 49.88 48.95 49.33
Supervised Uncertainty Quantification Method
SUQ 69.14 82.84 84.57 83.25
Table 3: AUC-PR for Yes-or-No tasks.
Appendix C Experimental Results
C.1 Yes-or-No Tasks
We only compare the uncertainty-based and SUQ methods on Yes-or-No tasks because the model’s answers exclusively contain “Yes” or “No”. The random baseline is calculated as the ratio of non-factual examples to the total number of examples. Our results in Table 3 demonstrate that SUQ method outperforms the uncertainty-based methods, exceeding AvgProb by about $4.5\%$. AvgProb achieves AUC-PR values nearly double those of the random baseline across all datasets and models, indicating it is an efficient method since it is an unsupervised method. In addition, the probability Prob overall performs better than the entropy Ent in Yes-or-No tasks.
Figure 6: Comparison of AvgProb with and without period in Yes-or-No tasks. V1.5-7b, V1.6-7b, and V1.6-Mis refer to LLaVA-v1.5-7b, LLaVA-v1.6-7b and LLaVA-v1.6-Mistral, respectively.
C.2 Impact of Unimportant Tokens for Uncertainty-based Methods
It is observed that the performance of uncertainty methods on the Open-ended tasks is not as good as on the Yes-or-No tasks. To explore the reasons, we append a period, resulting in “Yes.” or “No.” in the Yes-or-No task. In this case, probability $p$ and entropy $\mathcal{H}$ of two tokens need to be considered. Table 4 shows the detailed performance for Yes-or-No tasks with period using LLaVA-V1.6-Mistral. The AvgProb scores decline to 81.70 from 92.04 for adversarial, to 85.10 from 96.99 for random, to 81.74 from 95.06 for popular, and to 65.06 from 79.15 for GQA datasets, indicating the performance of AvgProb is severely affected by unimportant tokens. The other three metrics, AvgEnt, MaxProb, and MaxEnt, nearly equal to the baseline and play no role in hallucination detection. Figure 6 shows the comparison, indicating the performance of AvgProb is severely affected by unimportant tokens.
Method Mini LLaVA LLaVA LLaVA
GPT-v4 v1.5-7b v1.5-7b v1.6-Mistral
@@ C.2 Impact of Unimportant Tokens for Uncertainty-based Methods
C.3 Optimal Hidden Layers for Effective SUQ Method
@@
C.4 The comparison of clear and blurred images.
The comparison of clear and blurred images is illustrated in Figure 8.
@@
C.5 The performance of different methods in LLaVA-v1.6-mistral-7b.
Figure 9 visually presents the performance of different methods in LLaVA-v1.6-mistral-7b on IHAD.
Figure 9: PR-Curve of detecting non-factual sentences in LLaVA-v1.6-Mistral on IHAD.

### v2

@@ Abstract
*Equal contribution.†Corresponding author.
Reference-free Hallucination Detection for
Large Vision-Language Models
${\textbf{Qing Li}}^{1*}$, ${\textbf{Jiahui Geng}}^{1*}$, ${\textbf{Chenyang Lyu}}^{1}$, ${\textbf{Derui Zhu}}^{2}$, ${\textbf{Maxim Panov}}^{1}$, ${\textbf{Fakhri Karray}}^{1\dagger}$ 1 Mohamed bin Zayed University of Artificial Intelligence {qing.li, jiahui.geng, chenyang.lyu, maxim.panov, fakhri.karray}@mbzuai.ac.ae 2 Technical University of Munich derui.zhu@tum.de
@@ 1 Introduction
Large vision-language models (LVLMs) like LLaVA Liu et al. (2024), MiniGPT-4 Zhu et al. (2024b) have demonstrated remarkable capabilities in understanding and generating complex visual and textual content. However, an emerging concern with these models is their tendency towards hallucination Zhou et al. (2024); Zhu et al. (2024a); Geng et al. (2024b). For instance, a model might describe an object or event in the generated text that is not present in the input image. This phenomenon poses challenges for the reliability of LVLMs, highlighting intrinsic limitations of current models in maintaining coherence between visual inputs and textual descriptions Huang et al. (2023a).
Most existing methods of assessing hallucinations rely on external models. Li et al. (2023) propose POPE, which requires automated segmentation tools like SEEM Zou et al. (2024) to identify present and absent objects in images. Faithscore Jing et al. (2023) utilizes a visual entailment model Xie et al. (2019) to validate the factuality, focusing on object attributes and their relationships. Nonetheless, these approaches increase the computational cost and are restricted by the capabilities of these models.
@@
Reference-free methods are a promising approach to address this issue, primarily utilizing the model’s own knowledge. They have attracted great research interest in Large Language Models (LLMs), with typical methods including the uncertainty-based methods Hu et al. (2023); Geng et al. (2024a); Huang et al. (2023b); Vazhentsev et al. (2023); Fadeeva et al. (2023), consistency-based methods Manakul et al. (2023) and supervised uncertainty quantification (SUQ) approach Chen et al. (2023); Azaria and Mitchell (2023). However, their potential in the context of LVLMs is unknown. They have not yet been systematically studied, possibly due to the complexity of multimodal data. We are committed to bridging this gap.
In this work, we collect and implement a battery of state-of-the-art methods, including four metrics of uncertainty-based methods, four variants of consistency-based methods, and the supervised uncertainty quantification method. The experiments are conducted on five LVLMs with different types, versions, and sizes on Yes-and-No and Open-ended tasks. The main contributions of this paper are summarized as follows:
@@ 2 Reference-free Hallucination Detection Methods
Method Mini-GPT-4v LLaVA-v1.5-7b LLaVA-v1.6-vicuna-7b LLaVA-v1.6-mistral-7b
Adv Random 48.73 48.73 48.73 48.73
Uncertainty-based Methods
AvgProb 77.43 86.82 91.80 92.04
AvgEnt 48.89 50.05 48.90 48.92
Supervised Uncertainty Quantification Method
SUQ 87.23 92.35 93.32 93.50
Ran Baseline 48.73 48.73 48.73 48.73
Uncertainty-based Methods
AvgProb 86.62 95.89 96.23 96.99
AvgEnt 48.76 50.25 49.91 50.41
Supervised Uncertainty Quantification Method
SUQ 94.44 97.86 98.68 98.33
Pop Baseline 48.73 48.73 48.73 48.73
Uncertainty-based Methods
AvgProb 82.11 92.73 93.57 95.06
AvgEnt 48.50 49.76 50.63 51.39
Supervised Uncertainty Quantification Method
SUQ 91.02 97.85 98.11 98.50
GQA Baseline 48.47 48.47 48.47 48.47
Uncertainty-based Methods
AvgProb 66.87 78.45 80.31 79.15
AvgEnt 48.75 49.88 48.95 49.33
Supervised Uncertainty Quantification Method
SUQ 69.14 82.84 84.57 83.25
Table 1: AUC-PR for Yes-or-No tasks.
Uncertainty-based methods have the hypothesis that when LVLMs are uncertain about generated information, generated tokens often have higher uncertainty, shown in Figure 1(b). Uncertainty-based methods are extensively studied for hallucination detection in LLMs Hu et al. (2023); Geng et al. (2024a); Huang et al. (2023b); Fadeeva et al. (2024), and are also used as benchmark techniques Zhu et al. (2024a); Manakul et al. (2023). We adopt the four proposed metrics: AvgProb, AvgEnt, MaxProb, MaxEnt to aggregate token-level uncertainty to measure sentence-level and passage-level uncertainty. Appendix A.1 provides detailed descriptions of these metrics.
@@ 3 Study Design
Models. We evaluate five representative categories of LVLMs: MiniGPT-4v Zhu et al. (2024b), LLaVA-v1.5-7b Liu et al. (2023), LLaVA-v1.6-vicuna-7b, LLaVA-v1.6-mistral-7b Liu et al. (2024) and LLaVA-v1.6-vicuna-13b. These models, featuring different architectures and strong performance, are used as benchmarks in various works Huang et al. (2023a); Zhou et al. (2024). We only test models of a 7-billion and 13-billion size due to the limited computational resources.
@@
4 Experimental Results and Analysis
The comprehensive results on Open-ended tasks are presented in Tables 1,  2 and 3. The numbers in the pink and yellow background, respectively, represent the best results for the uncertainty-based methods and consistency-based methods, while the numbers in bold indicate the overall best results. We utilize AUC-PR Davis and Goadrich (2006), which stands for the area under the precision-recall curve, to objectively evaluate the effectiveness of each model. The higher the value of AUC-PR, the stronger the ability of this method for hallucination detection.
@@ 4.1 Effectiveness of Various Methods
Hallucination Detection on Yes-or-No Tasks. We only compare the uncertainty-based and SUQ methods on Yes-or-No tasks because the model’s answers exclusively contain “Yes” or “No”. The random baseline is calculated as the ratio of non-factual examples to the total number of examples. Our results in Table 1 demonstrate that SUQ method outperforms the uncertainty-based methods, exceeding AvgProb by about $4.5\%$. AvgProb achieves AUC-PR values nearly double those of the random baseline across all datasets and models, indicating it is an efficient method since it is an unsupervised method. In addition, the probability Prob overall performs better than the entropy Ent in Yes-or-No tasks.
@@
4.2 Robustness and Limitations of SUQ
@@
4.3 Impact of Image Clarity
@@
4.4 Impact of Data Source
@@
Others. We have also explored the impact of unimportant tokens for uncertainty-based methods and the optimal hidden layers for SUQ. They are discussed respectively in Appendices C.1 and C.2.
@@ 5 Conclusion
This paper systematically studies reference-free approaches for detecting the hallucination of LVLMs. Our results indicate that the SUQ method performs best as a supervised method, but it relies on sufficient training data, which could limit its applicability in certain scenarios. Consistency-based methods outperform uncertainty-based methods; however, they require more computation and are more suitable for black-box models. Uncertainty-based methods are particularly well-suited for tasks similar to Yes-or-No questions.
@@
Three major limitations are identified in this work. First, we did not compare with reference-based methods. The performance of reference-based methods depends on the performance of external models and is specific to certain tasks. In this article, we primarily aim to explore the potential of reference-free methods, which have a broader range of applications. Second, we did not classify types of hallucinations; limited by resources, even with fine-grained annotations, we only marked whether sentences contained hallucinations without distinguishing types such as object existence errors, attribute errors, text recognition errors, and fact-conflicting errors. This might be where SUQ methods have an advantage. Third, we only studied white-box models. Recent research indicates that white-box models can be proxies for black-box models to calculate their confidence regarding specific issues. Unigram method is one such proxy model approach. In the future, we could use more complex models, closer in architecture to the target models, as proxies to explore their potential in hallucination detection.
@@ B.1 Dataset
GPT-v4 v1.5-7b v1.6-vicuna-7b v1.6-mistral-7b
@@
B.2 Experimental Setup
Appendix C Experimental Results
C.1 Impact of Unimportant Tokens for Uncertainty-based Methods
It is observed that the performance of uncertainty methods on the Open-ended tasks is not as good as on the Yes-or-No tasks. To explore the reasons, we append a period, resulting in “Yes.” or “No.” in the Yes-or-No task. In this case, probability $p$ and entropy $\mathcal{H}$ of two tokens need to be considered. Table 4 shows the detailed performance for Yes-or-No tasks with period using different methods. Specifically, for LLaVA-v1.6-mistral, the AvgProb scores decline to 81.70 from 92.04 for Adv, to 85.10 from 96.99 for Ran, to 81.74 from 95.06 for Pop, and to 65.06 from 79.15 for GQA datasets, indicating the performance of AvgProb is severely affected by unimportant tokens. The other three metrics, AvgEnt, MaxProb, and MaxEnt, nearly equal to the baseline and play no role in hallucination detection.
C.2 Optimal Hidden Layers for Effective SUQ Method
@@
C.3 The comparison of clear and blurred images.
The comparison of clear and blurred images used in 4.3 is illustrated in Figure 7.